# ExoPredicator: Learning Abstract Models of Dynamic Worlds for Robot Planning

### v1

@@ 7 Related Works
Learning Abstractions for Planning. Early work learned STRIPS or NDR transition rules from demonstrations given a fixed predicate set (Pasula et al., 2007; Silver et al., 2021; 2022; Chitnis et al., 2022). Recent methods invent new predicates to improve generalisation (Silver et al., 2023; Liang et al., 2024); however, all assume instantaneous deterministic effects. ExoPredicator extends predicate‑invention to environments with exogenous dynamics and delayed causal effect.
@@ 8 Conclusion
We presented ExoPredicator, an integrated approach for learning and planning with causal processes in environments with exogenous dynamics and delayed effects. Our method demonstrates the ability to learn abstract world models from limited data, generalizing to new tasks with unseen objects and goals across various simulated environments, and outperforming key baselines. Future work will scale the framework to more complex, larger-scale environments, enhance learning with foundation models, and explore the interplay between skill and world modeling.
@@ A.4 Process Learning details
Given a set of trajectories $\mathcal{D}$, a set of predicates $\Psi$, and the agent’s known endogenous processes $\mathcal{L}_{en}$, we learn the set of exogenous processes $\mathcal{L}_{ex}$. Our method follows Chitnis et al. (2022) in assuming that for any given effect (a unique pair of add/delete atoms), there is at most one exogenous process that causes it. While this prevents learning multiple distinct causes for the same outcome, it significantly simplifies the search problem. Any lost expressivity can be recovered by inventing more nuanced predicates. The learning algorithm proceeds in five steps:
@@
    Cluster. We then cluster the remaining segments based on their effects. We assume that each exogenous process has a single, atomic effect (e.g., one predicate changing from false to true). [^2] This imposes no loss of generality, because separate processes can be learned for each changed predicate. If a segment involves multiple predicate changes, we duplicate it into multiple clusters—one for each change—allowing us to learn a separate process for each atomic effect.
@@
    $L_{C^{*},E}=\operatorname*{arg\,max}_{L_{C_{i},E}}\log p(L_{C_{i},E}|\mathcal{D})=\operatorname*{arg\,max}_{L_{C_{i},E}}\left(\log p(\mathcal{D}|L_{C_{i},E})+\log p(L_{C_{i},E})\right)$
    where $\log p(\mathcal{D}|L_{C_{i},E})$ is the approximate marginal likelihood from the previous section and $\log p(L_{C_{i},E})$ is a minimum description length prior that penalizes overly complex conditions.
@@ C.1.5 Fan
C.2 Further Planning Statistics

### v3

@@ 7 Related Works
Hierarchical Reinforcement Learning. HRL uses temporally extended actions to address long-horizon decision-making (Barto & Mahadevan, 2003). While many skill-learning approaches exist, they typically adopt the (semi-)Markov assumption at the option level: the distribution over outcomes depends only on the initiation state and the chosen option (Masson et al., 2016; Nasiriany et al., 2022; Mishra et al., 2023). In such formulations, environment-driven or exogenous dynamics can be absorbed into the option’s transition model, but are not usually represented as separate causal processes with their own activation conditions and delay distributions. Our approach is complementary: we learn an explicit model for these external dynamics and stochastic delays, allowing a fixed set of skills to be used more flexibly as the environment evolves and enabling more generalizable planning.
@@
Learning Abstractions for Planning. Early work learned STRIPS or NDR transition rules from demonstrations given a fixed predicate set (Pasula et al., 2007; Silver et al., 2021; 2022; Chitnis et al., 2022). Recent methods invent new predicates to improve generalisation (Silver et al., 2023; Liang et al., 2024); however, they typically abstract dynamics through agent-centric actions, without explicitly modeling exogenous causal processes that unfold over time in the background,. ExoPredicator extends predicate‑invention to environments with exogenous dynamics and delayed causal effect.
@@ 8 Conclusion
We presented ExoPredicator, an integrated approach for learning and planning with causal processes in environments with exogenous dynamics and delayed effects. Our method demonstrates the ability to learn abstract world models from limited data, generalizing to new tasks with unseen objects and goals across various simulated environments, and outperforming key baselines. Future work will scale the framework to more complex, noisier, larger-scale environments, enhance learning with foundation models, and explore the interplay between skill and world modeling.
@@ A.3 Probabilistic Model and ELBO Derivation
For experiments, we use a categorical variational distribution over the discrete support of the delays. The variational parameters are initialized to be uniform over this support.
@@ A.4 Process Learning details
Given a set of trajectories $\mathcal{D}_{\text{abs}}$, a set of predicates $\Psi$, and the agent’s known endogenous processes $\mathcal{L}_{en}$, we learn the set of exogenous processes $\mathcal{L}_{ex}$. Our method follows Chitnis et al. (2022) in assuming that for any given effect (a unique pair of add/delete atoms), there is at most one exogenous process that causes it. While this prevents learning multiple distinct causes for the same outcome, it significantly simplifies the search problem. Any lost expressivity can be recovered by inventing more nuanced predicates. The learning algorithm proceeds in five steps:
@@
    Cluster. We then cluster the remaining segments based on their effects. We assume that each exogenous process has a single, atomic effect (e.g., one predicate changing from false to true). [^2] This imposes no loss of generality, because separate processes can be learned for each changed predicate. If a segment involves multiple predicate changes, we duplicate it into multiple clusters—one for each change—allowing us to learn a separate process for each atomic effect. This induces a deterministic partition of segments by effect, so the step introduces no clustering hyperparameters.
@@
    $L_{C^{*},E}=\operatorname*{arg\,max}_{L_{C_{i},E}}\log p(L_{C_{i},E}|\mathcal{D}_{\text{abs}})=\operatorname*{arg\,max}_{L_{C_{i},E}}\left(\log p(\mathcal{D}_{\text{abs}}|L_{C_{i},E})+\log p(L_{C_{i},E})\right)$
    where $\log p(\mathcal{D}_{\text{abs}}|L_{C_{i},E})$ is the approximate marginal likelihood from the previous section and $\log p(L_{C_{i},E})$ is a minimum description length prior that penalizes overly complex conditions.
@@ C.1.5 Fan
C.2 Further Learning and Planning Statistics
Each online learning iteration of ExoPredicator takes approximately 30–360 seconds wall-clock time depending on the domain, using 32 CPU cores and no GPU. In comparison, the HRL baseline requires on average about 600 seconds per online iteration, and its runtime increases as it accumulates more data (taking longer to converge), reaching up to  3200 seconds in some cases. The VLM planning baseline does not perform online learning, but can incur substantial test-time cost because it issues computationally expensive VLM calls every time it generates a plan.
Monetaryly, assuming Gemini 2.5 Pro with standard pricing, a representative online iteration consuming a combined 10,000 input tokens and 1,000 output tokens would cost approximately \$0.0225 in total: \$0.0125 for inputs (10,000 / 1,000,000 $\times$ 1.25) and \$0.01 for outputs (1,000 / 1,000,000 $\times$ 10).
The following table lists the success rate and planning time statistics.
@@ C.2 Further Learning and Planning Statistics
C.3 Priors and Distributions
In all experiments, we model process delays with a truncated discrete Gaussian distribution over positive integers $\{1,\dots,300\}$. The distribution is parameterized by a log-mean and a log-standard-deviation.
The log-mean, log-standard-deviation, and log-process-weights are initialized from a normal distribution with mean 0 and standard deviation $0.01$.